OpenStudy (anonymous):

y = 7/x^2 + 10 Find f(x) and g(x) so the function can be expressed as y = f(g(x)).

Parth (parthkohli):

Okay, so you can introduce some other stuff here. Let's figure it out. To start with, we are dividing 1 by a number's square + 10 and multiplying 7 to it... right?

Parth (parthkohli):

So the first thing you are doing can be said as the inner function. So,\[ f(x) = {1 \over x^2 +10}\]Now what do you think is the outer function?

OpenStudy (anonymous):

g(x)=7/1

OpenStudy (anonymous):

i don't think that'll work... check to see if you get the correct expression if you compute f(g(x))...

OpenStudy (anonymous):

\[f(x)=(\frac{ 1 }{ x^2 } +10) g(x)=\frac{ 7 }{ 1 }\] \[f(g)(x)=\frac{ 7 }{ x^2 } +10\] I think, i really don't get this so i'm kinda trying

OpenStudy (anonymous):

wait.. is the original expression this???: \(\large y=\frac{7}{x^2+10} \) ???? or is it something else?

OpenStudy (anonymous):

the original form is \[\frac{ 7 }{ x^2 } +10\]